Jump to main content or area navigation.

Contact Us

Waste Site Cleanup & Reuse in New England

Atlas Tack Corporation Superfund Site
Bi-Weekly Remedial Activities Update
Report # 20: November 27 - December 1, 2006

« Back to Atlas Tack Superfund Site Fact Sheet

1. ACTIVITIES PERFORMED DURING THE WEEK OF NOVEMBER 27 - DECEMBER 1

  • Continued to Transport and disposal (T&D) of Subtitle D soils from the Commercial and Debris Area (CID) and the East Fill Area
  • T&D Subtitle-D oversize soils
  • Continue to T&D COMM-97-soils to Marion landfill from stockpiles CD-25 and WF-03
  • Continued to excavate soils in the NW-Fill Area; grids: J4, K4, L4, M4, L3, K3
  • Removed sections of the interior fence and vegetation east of Commercial Area
  • Provide topographic survey of the marsh area south of the Hurricane Barrier
  • Return to the western boundary of the CID and excavate cells A4, C2, D2 and G1

Ongoing Site Activities:

  • Collected waste characterization soil samples from soils originating from NW-Fill Area for off site landfill profiling and disposal
  • Collected confirmatory soil samples from the floors of the NW-Fill Area (all floor samples collected to date pass clean-up criteria
  • Continue water handling activities
  • Continued to inspect trucks prior to exiting the site
  • Continued to provide flaggers to cross trucks across the bike paths along Adams and Tripp Streets
  • Continued stockpile management activities
  • Continue with air monitoring program

Starting November 27, 2006, weekly updates changed to a biweekly update (every two weeks). This may change back to a weekly schedule if EPA warrants this.

2. PROJECTED ACTIVITIES FOR THE WEEK OF DECEMBER 4 - DECEMBER 8 AND DECEMBER 11 - DECEMBER 15

  • Begin excavation in the Lagoon Area.
  • Continued to T&D Subtitle D soils from stockpiles East and Northwest Fill Areas
  • Continue to T&D COMM-97 soils to Crapo landfill
  • Continue with excavation activities in the Lagoon Area
  • Collect waste characterization soil samples from soils originating from Lagoon Area

Ongoing Site Activities:

  • Continue to collect pre and post excavation survey data of the excavations
  • Continue with stockpile management activities
  • Continue water handling activities
  • Continued to inspect trucks prior to exiting the site
  • Continued to provide flaggers to cross trucks across the bike paths along Adams and Tripp Streets
  • Continue with air monitoring program

3. AIR MONITORING INFORMATION

WESTON monitored for particulate (dust) during intrusive activities performed during this reporting period. WESTON recognizes EPA's level of concern for "PM10" to caution sensitive groups, including children, the elderly, and people with heart or lung disease, about exposure to coarse dust particles. The PM10 level of concern for sensitive groups is 150 µg/m3 for 24-hour average exposure.

Perimeter Air Monitoring is performed at up to (6) locations throughout a given work day. Readings are collected at 60 second intervals throughout the day and reported as a daily average.

If exceedances are identified, an increased effort of engineering controls (water suppression) will be utilized to adhere to the PM10 level of concern. If engineering controls appear to be inadequate, any dust generating activities will stop. Current dust control practices will be re-evaluated until concentrations are reduced to below the PM10 standard.

Results for 16 November – 28 November 2006 can be reviewed in the table shown below.

Weekly Dust Monitoring Summary
Week #20: Results for Thursday through Wednesday**

The EPA's PM10 level of concern for sensitive groups is 150 µg/m3 for 24-hour average exposure

Date Pleasant Street (µg/m3) Adams Street (µg/m3) Tripp Street (µg/m3) Egypt Road (µg/m3) Church Street (µg/m3)
11/16/2006 11 0 0 NSC NSC
11/17/2006 Precip Precip Precip Precip Precip
11/20/2006 0 0 0   0
11/21/2006 0 0 0 Precip 0
11/22/2006 0 0 0   0
11/23/2006 Holiday Holiday Holiday Holiday Holiday
11/24/2006 Holiday Holiday Holiday Holiday Holiday
11/27/2006 0 0 0   0
11/28/2006 0 0 0   0
11/29/2006          

Notes:
All readings reflect the daily average calculated over the course of the workday.
All readings collected with a Thermo Electron PDR 1000.
NSC: No sampling conducted; No intrusive work conducted in that area.
Precip: No sampling conducted due to precipitation.
*: Monitoring conducted over 24-hour period.
**: No work and air monitoring occurs on weekend days

4. T&D "BI-WEEKLY" SUMMARY

DATE
DESTINATION
WASTE DETERMINATION
CY
TONS
STOCKPILES
28 Nov. 06 TREE Subtitle-D     EA-17
27 Nov. 06 TREE Subtitle-D     EA-19 and EA-5
22 Nov. 06 TREE Subtitle-D 432.83 649.25 EA-19 and EA-16
21 Nov. 06 TREE Subtitle-D 358.78 538.17 EA-16
20 Nov. 06 TREE Subtitle-D 268.44 402.66 CD-17, 18, 20 and 22
17 Nov. 06 MARION COMM-97 344.37 516.55 CD-24
17 Nov. 06 BOURNE COMM-97 Oversize 252.64 303.17 EA-15, EA-27, EA-21
16 Nov. 06 BOURNE COMM-97 Oversize 333.48 400.18 EA-15, EA-27, EA-21

5. T&D TOTAL SUMMARY

WK/E
DESTINATION
WASTE DETERMINATION
ORIGIN
CY
TONS
1 Dec. 06 TREE Sub-D Oversize EAST FILL 2,500 3,750
24 Nov. 06 TREE Sub-D Oversize CID/ EAST FILL 1,060 1,590
17 Nov. 06 Bourne Sub-D Oversize CID/ EAST FILL 1,981 2,481
10 Nov. 06 TREE SUB-D CID 1,130.37 1,695.56
03 Nov. 06 Marion COMM-97 CID/EAST FILL 1,881.51 2,822.27
27 Oct. 06 T&D DID NOT OCCUR THIS WEEK
20 Oct. 06 T&D DID NOT OCCUR THIS WEEK
13 Oct. 06 T&D DID NOT OCCUR THIS WEEK
06 Oct. 06 Marion COMM-97 CID 843.04 1,264.56
29 Sept. 06 Marion COMM-97 CID 449.63 674.45
T&D TOTAL
10,670.45 15,525.70

6. EXCAVATION SUMMARY

WEEKS /E 11/24/06 – 12/01/06
TO DATE
ACTIVITIES INITIAL TOTAL QUANTITY INITIAL TOTAL QUANTITY
Commercial Area Excavation None 2,655 CY
C.I.D. Excavation None 7,083 CY
East Fill Excavation Complete 9,046 CY
NW-Fill Area 1,400 CY 7,800 CY
 

Jump to main content.